# 3DGSR: Implicit Surface Reconstruction with 3D Gaussian Splatting

### v1

@@ 4.1. Connect 3D Gaussians with Implicit SDF
The vanilla 3DGS method optimizes millions of Gaussians to accurately fit the training views, enabling photorealistic synthesis of novel views. However, its main emphasis is on fitting the target views, which does not guarantee high-quality 3D geometries. Moreover, the 3D Gaussians should be constrained by the underlying geometry (i.e., surfaces) of the 3D scene, which is absent in current methods. Consequently, 3DGS faces limitations in accurately reconstructing 3D surfaces (see Fig. 5).
@@
Here, we introduce neural implicit SDF, denoted as $f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{1}$, which maps each 3D position $x\in\mathbb{R}^{3}$ to its signed distance from the target surface. This is implemented using multi-resolution hash grids (Müller et al., 2022) in conjunction with a single-layer MLP. The surface $\mathrm{S}$ of a scene can then be derived by using the zero-level set of its SDF, defined by:
(6) $\mathrm{S}=\{x\in\mathbb{R}^{3}\>|\>f(\mathbf{x})=0\}.$
@@
Next, to link the SDF with 3D Gaussians, we introduce a differentiable SDF-to-opacity transformation function. This function should adhere to the constraint that points closer to the surface contribute more to rendering a pixel and, therefore, have a higher opacity value. Consequently, we select a bell-shaped function $\Phi_{\beta}(f(x))$ controlled by a hyperparameter $\gamma$, expressed as:
@@
where $\beta$ is a learnable parameter controlling the shape of the function.
This function transforms SDF values at a query location into the opacity of the corresponding Gaussian at that location. By linking the SDF with 3D Gaussians using a differentiable function, the SDF field can be updated through the optimization of 3D Gaussians by minimizing the photometric loss:
@@
This loss function encourages the Gaussian at a location $x$ that contributes more to rendering a pixel to have a higher opacity value $\Phi_{\beta}(f(x))$. This, in turn, encourages the SDF $f(x)$ to be closer to 00 and, therefore, nearer to the surface. The high-quality rendering capability of 3DGS allows for learning detailed image structures, which facilitates the reconstruction of detailed structures (see Fig. 5).
In addition to $\mathcal{L}_{c}$, we also impose a constraint as Eq. (9) to maintain consistency between the rendered normal $\mathcal{N}$, obtained directly from Gaussians via Eq. 5, and a pseudo normal $\hat{\mathcal{N}}$, computed from the rendered depth $\mathcal{D}$ derived from 3D Gaussians under the local planarity assumption.
(9) $\mathcal{L}_{n}=\|\mathcal{N}-\hat{\mathcal{N}}\|_{2}.$
Further, to encourage $f$ to approximate a valid signed distance function, we follow the Eikonal equation (Gropp et al., 2020) to penalize the deviation of $f$:
(10) $\mathcal{L}_{\text{eik}}=\mathbb{E}_{x}[(\|\triangledown f-1\|)^{2}].$
Additionally, we enforce that the distribution of Gaussians adheres to the surface by constraining the SDF values at their centers to be close to zero:
(11) $\mathcal{L}_{\text{pt}}=\|f(x_{g})\|_{1},$
where $x_{g}$ denotes the centers of 3D Gaussians.
@@ 4.2. Regularization with Volumetric Rendering
By integrating the SDF field with 3D Gaussians and training them jointly, as discussed in Section 4.1, we can attain high-fidelity rendering results with an optimized SDF field. However, due to the sparse and discrete nature of 3D Gaussians, the aforementioned supervision signals are only available for locations occupied by 3D Gaussians, predominantly near the surface. This is insufficient to regularize the continuous SDF, resulting in artifacts at locations not covered by Gaussians, such as floating artifacts in unoccupied free spaces, as illustrated in Fig. 3.
Figure 3. Comparison of our method with and without volumetric rendering regularization. The incorporation of volumetric rendering regularization proves to be highly effective in eliminating floating artifacts in free space where Gaussians are absent.
Figure 4. Comparison of volume rendering results and Gaussian splatting results on the TNT dataset. The resemblance demonstrates the efficacy of our learning process in aligning the rendered geometric attributes - depth and normal - derived from volumetric rendering (which is based on the learned SDF) and those obtained from 3DGS.
Since volumetric rendering employs ray-casting to render a pixel, it samples points along a ray that includes both occupied and free points. We propose incorporating volumetric rendering for optimization. The depth map and normal map of volume rendering can effectively represent the occupancy distribution in the whole space, while those of sparse Gaussians aligns well with the observed image. This inspires us to adopt the rendering results of Gaussians’ geometry attributes as the guidance of volume rendering, which can be utilized as the medium for propagating gradients. Specifically, we employ volumetric rendering to render depth (normals) and supervise them using depth (normals) derived from Gaussians using Eq. (5). This can provide regularizations to SDF values at locations uncovered by Gaussians.
@@
where $\delta^{r}_{i}$ is the distance between adjacent sample points and $\phi_{s}$ is the Sigmoid function.
@@
where $\mathcal{R}$ is the union of all training rays. These constraints regularize all the sampling points along the training rays, resulting in dense supervision for the implicit field and eliminating floating artifacts. As shown in Fig. 2, there exist noisy SDF distributions (negative value) floating over the free space between the camera and the object, which is then removed by applying the regularization with volumetric rendering. Accurate geometry can be obtained through our careful training strategy and model design. We visualize the optimized depth map and normal map of both volume rendering and Gaussian rendering in Fig 4.
The overall training objective is defined as a combination of all aforementioned losses:
(16) $\mathcal{L}=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{n}+\lambda_{2}\mathcal{L}_{\text{eik}}+\lambda_{3}\mathcal{L}_{\text{pt}}+\lambda_{4}\mathcal{L}_{\text{vd}}+\lambda_{5}\mathcal{L}_{\text{vn}},$
where $\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5}$ are hyperparameters that control the relative importance of each loss term.
@@ 5.1. Experimental Settings
Our evaluation encompasses three datasets: the NeRF synthetic dataset (Mildenhall et al., 2021), the DTU real-captured dataset (Jensen et al., 2014), and the Tanks&Temples (TNT) large-scale dataset (Knapitsch et al., 2017). The NeRF synthetic dataset consists of eight objects, each rendered via path tracing to generate 100 views. For testing, we select an eighth of these images, allocating the remainder for training. The DTU and TNT datasets are processed for camera poses as per (Gao et al., 2023), with image and mask processing following the methodology outlined in (Liu et al., 2020a). The image resolutions for the DTU and TNT datasets are $1600\times 1200$ and $1920\times 1080$ pixels, respectively.
Comparisons
We evaluated our method from both surface reconstruction quality and rendering quality of novel-view synthesis. We benchmark our algorithm against several state-of-the-art methods, including NeRF (Mildenhall et al., 2021), NeuS (Wang et al., 2021), NeuSG (Chen et al., 2023a), RelighableGaussian (Gao et al., 2023), and NeRF2Mesh (Tang et al., 2023b), Neuralangelo (Li et al., 2023b), PET-NeuS (Wang et al., 2023b), BakedSDF (Yariv et al., 2023), VolSDF (Yariv et al., 2021). The quality of image synthesis is quantified using the Peak Signal-to-Noise Ratio (PSNR). For surface reconstruction accuracy, we employ the Chamfer distance and the F1 score as our metrics (Knapitsch et al., 2017). Furthermore, we also compare the rendering quality of our method with some works that solely focus on novel view synthesis tasks (Tri-MipRF (Hu et al., 2023), Zip-NeRF (Barron et al., 2023), 3DGS (Kerbl et al., 2023)).
@@
Ours 34.85 26.08 35.17 36.88 34.90 30.03 36.44 31.48 33.23
@@
Ours 1.01 0.95 0.69 1.92 1.35 1.35 1.15 3.35 1.50
@@
Ours 0.68 0.84 0.70 0.39 1.16 0.87 0.77 1.48 1.26 0.87 0.69 0.80 0.42 0.64 0.60 0.81
@@
Table 2. Quantitative assessment on the DTU dataset with each cell colored to indicate the best and second best. Our method achieves the highest average quality of surface reconstruction and the lowest Chamfer-$\mathcal{L}_{1}$ distance with real-world data.
Figure 5. Qualitative comparisons of reconstructed results between NeRF2Mesh (Tang et al., 2023b), Relightable 3D Gaussian (Gao et al., 2023), and our method on the NeRF Synthetic dataset. Our method yields the highest-quality surface reconstruction preserving fine details and intricate structures.
@@
Figure 6. Qualitative comparisons of reconstructed results between VolSDF (Yariv et al., 2021), Relightable 3D Gaussian (Gao et al., 2023), and our method on the DTU dataset. Our method yields the highest-quality surface reconstruction preserving fine details and intricate structures.
Implementation Details
Our work is implemented by PyTorch and CUDA and optimized by AdamW optimizer. We bind the Gaussian points to the SDF field for the whole training process to ensure that the Gaussian rendering process can provide a coarse SDF field, that can boost the converging speed of volumetric rendering. To ensure stable optimization, we introduce the warm-up strategy to make sure the dynamic Gaussian points can move around to the surface and provide an accurate supervised signal. The warm-up steps are set as 3000 for the Nerf synthetic dataset and 5000 for the TNT dataset. During the volumetric rendering stage, we sample $N=1024$ rays per step.
For the synthetic data, we initialize the implicit function as a sphere, and the position of Gaussian is determined by the zero-level surface of the SDF field. For the real-world scenes, which frequently exhibit intricate geometry and noised pose, we adopt the SFM points as the initialized Gaussian position, and the network is still initialized as a sphere. Furthermore, we also apply the sparse MVS points to ensure the consistency between the rendered depth $\mathcal{D}$ and projected depth map $D_{mvs}$ with $\mathcal{L}_{1}$ loss. Our models are trained for 30000 iterations, and all the experiments are conducted on a single NVIDIA GeForce RTX 3090 GPU.
5.2. Ablation Study

### v2

@@ 4.1. Alignment of Surfaces and Gaussians
The vanilla 3DGS method optimizes millions of Gaussians to accurately fit the training views, enabling photorealistic synthesis of novel views. However, its main emphasis is on fitting the target views, which does not guarantee high-quality 3D geometries. Moreover, the 3D Gaussians should be constrained by the underlying geometry (i.e., surfaces) of the 3D scene, which is absent in existing methods.
@@
Here, we introduce neural implicit SDF, denoted as $f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{1}$ for surface modeling, which maps each 3D position $x\in\mathbb{R}^{3}$ to its signed distance from the surface. This is implemented using multi-resolution hash grids (Müller et al., 2022) in conjunction with a single-layer MLP. The surface $\mathrm{S}$ of a scene can then be derived by using the zero-level set of its SDF, defined by:
(6) $\mathrm{S}=\{x\in\mathbb{R}^{3}\>|\>f(x)=0\}.$
Given the SDF $f$, the next question is to connect 3DGS with the implicit SDF to enable unified surface and appearance modeling and optimization. Here, we explore two coupling strategies: tight coupling with a differentiable SDF-to-opacity transformation function and loose coupling with distance and orientation regularization.
Tight Coupling Strategy.
@@
Intuitively, points near the surface will have large opacity values and contribute more to rendering images. Thus, we introduce a differentiable SDF-to-opacity transformation function to tightly couple the SDF field and 3DGS. Consequently, we select a bell-shaped function $\Phi_{\beta}(f(x))$ controlled by a hyperparameter $\beta$, expressed as:
@@
where $\beta$ is a learnable parameter controlling the shape of the function. This function transforms SDF values at a query location into the opacity of the corresponding Gaussian at that location. By linking the SDF with 3D Gaussians using a differentiable function, the SDF field can be updated through the optimization of 3D Gaussians by minimizing the photometric loss:
@@
This loss function encourages the Gaussian at a location $x$ that contributes more to rendering a pixel to have a higher opacity value $\Phi_{\beta}(f(x))$. This, in turn, encourages the SDF $f(x)$ to be closer to 00 and, therefore, nearer to the surface.
Loose Coupling.
Although the tight coupling strategy can align Gaussian points with the implicit surface and enable unified optimization for enhanced reconstructions, it may slightly compromise rendering quality and produce unsatisfactory reconstruction results with visible artifacts (See Fig. 3). Upon investigation, we find that the surface and color appearances indeed exhibit different levels of detail and frequencies, which could be a contributing factor.
For example, as shown in Fig. 3, a chair with flat low-frequency surface areas may require numerous non-uniformly distributed Gaussian points to represent its high-frequency textures. In areas with dense Gaussian points, the tight coupling strategy enforces the SDF values to be closer to zero, potentially leading to noisy surface structures with unnecessary details (see Fig. 10) if the Gaussian locations are not accurate. This suggests that a tight alignment might not be the optimal solution.
Thus, we introduce a loose coupling strategy as an alternative to the tight coupling approach. The central concept involves using the surface derived from the SDF to regularize the distribution of 3D Gaussians. This strategy ensures that 3D Gaussians are positioned near the surface and their orientations are aligned with the surface normals. By aligning the surface normals instead of directly constraining the opacity, we can prevent interference with the supervision needed for high-frequency appearances.
Specifically, the direction of a 3D Gaussian at location $x_{g}$ is determined by the direction of its shortest axis of the covariance matrix, represented by $n_{g}$. The normal at $x_{g}$ in the SDF field is derived using $\nabla{f}(x_{g})$. Note we normalize $\nabla{f}(x_{g})$ to a unit vector. We then align the direction of the 3D Gaussian and the surface normal by minimizing the direction-aligned loss, which is defined as follows:
(9) $\mathcal{L}_{a}=||1-n_{g}\nabla{f}(x_{g})||.$
Additionally, we constrain the 3D Gaussians to be near the surface. To achieve this, we query the SDF value of each Gaussian point $x_{g}$ and calculate its nearest surface location using the gradient of the SDF field as follows:
(10) $x_{\text{nearest}}=x_{g}+f(x_{g})\cdot\nabla{f}(x_{g}),$
where $x_{\text{nearest}}\in R$ represent the nearest 3D points of point $x_{g}$ on the surface from SDF. Then, we can use $\mathcal{L}_{1}$ loss to minimize their distance. As shown in Fig. 3, despite its simplicity, the loose coupling strategy yields detailed and accurate surface reconstruction results where the 3D Gaussians near the surface with their orientations aligned with surface normals. Moreover, this approach also helps reduce the number of Gaussians by preventing the generation of Gaussian points that are far away from the surface. Our experiments demonstrate that we can decrease the number of Gaussian points by 20% without compromising, or even enhancing, the rendering quality. If not explicitly specified, we use the loose coupling strategy for alignment by default.
@@ 4.2. Regularization with Volumetric Rendering
By integrating the SDF field with 3D Gaussians and jointly training them, as discussed in Section 4.1, we can achieve high-fidelity rendering results with an optimized SDF field. However, due to the sparse and discrete nature of 3D Gaussians, the supervision signals mentioned earlier are only available for locations occupied by 3D Gaussians, primarily near the surface. This limited availability is insufficient to regularize the continuous SDF, leading to artifacts in locations not covered by 3D Gaussians, such as floating artifacts in unoccupied free spaces, as shown in Fig. 4.
Volumetric rendering employs ray-casting to render a pixel, sampling points along a ray that includes both occupied and free points. Thus, we propose incorporating volumetric rendering to optimize the dense SDF field. Specifically, we use volumetric rendering to render depth and normal, which can be directly derived from the SDF field, and align them with those derived from 3DGS for optimization. There are two main reasons for using rendered depth and normal instead of color. First, rendering color would require an additional color prediction MLP network, significantly increasing the learning time. Second, and more importantly, learning the SDF directly from geometric clues (e.g., depth and normal) can alleviate the burden of appearance modeling, accelerate the learning process, and enhance the quality.
@@
where $\phi_{s}$ is the Sigmoid function.
@@
where $\mathcal{R}$ is the union of all training rays. These constraints regularize all the sampling points along the training rays, resulting in dense supervision for the implicit field and eliminating floating artifacts. As shown in Fig. 4, there exist noisy SDF distributions (negative value) floating over the free space between the camera and the object, which is then removed by applying the regularization with volumetric rendering. Accurate geometry can be obtained through our careful training strategy and model design.
4.3. Training Objectives and Implementation Details
To train 3DGRS, our objective is to get high-quality rendering and reconstruction results. Thus, the final loss function is defined by
$\mathcal{L}=\mathcal{L}_{c}+\alpha_{1}\mathcal{L}_{vd}+\alpha_{2}\mathcal{L}_{vn}+\alpha_{3}\mathcal{L}_{a}+\alpha_{4}\mathcal{L}_{eik}$
where $\mathcal{L}_{eik}$ is the Eikonal equation (Gropp et al., 2020) to penalize the deviation of the magnitude of grad $f$ from unit length:
(15) $\mathcal{L}_{\text{eik}}=\mathbb{E}_{x}[(\|\triangledown f-1\|)^{2}].$
In our experiments, we set $\alpha_{1}=1.0,\alpha_{2}=0.1,\alpha_{3}=0.0001,\alpha_{4}=0.1$.
We implement the 3DGSR with CUDA and tinycudann (Müller, 2021), building upon the framework of 3DGS. We customize the normal and depth maps CUDA rendering kernel. During training, we increase the number of 3D Gaussian primitives following the adaptive control strategy in 3DGS. During the volumetric rendering stage, we sample $N=1024$ rays per step.
For the synthetic data, we initialize the implicit function as a sphere, and the position of Gaussian is randomly initialized. For the real-world scenes, which frequently exhibit intricate geometry and noisy poses, we adopt the SFM points as the initialized Gaussian position, and the network is still initialized as a sphere. Furthermore, we also apply the sparse SFM points to ensure the consistency between the rendered depth $\mathcal{D}$ and projected depth map $D_{sfm}$ with $\mathcal{L}_{2}$ loss. Specifically, when optimizing the pose using SFM algorithms like COLMAP, we can also obtain sparse 2D key points along with their corresponding 3D point locations. Given an image and its camera pose, we project visible 3D points into the camera space and use the z-value as the depth $D_{sfm}$ for the corresponding 2D key points. We then apply an $\mathcal{L}_{2}$ loss to enforce consistency between the rendered depth $\mathcal{D}$ and the projected sparse depth $D_{sfm}$ based on the measurements. Our models are trained for 30000 iterations, and all the experiments are conducted on a single NVIDIA RTX 3090 GPU.
@@ 5.1. Experimental Settings
Figure 5. Qualitative comparisons of different coupling methods and the influence of geometric constraint.
@@
2DGS (Huang et al., 2024) 35.05 26.05 35.57 37.36 35.10 29.74 35.09 30.60 33.07
Ours (Tight Coupling) 34.85 26.08 35.17 36.88 34.90 30.03 36.44 31.48 33.23
Ours (Loose Coupling) 35.69 26.35 35.65 37.98 36.19 30.31 37.12 31.65 33.86
@@
Ours (Tight Coupling) 1.01 0.95 0.69 1.92 1.35 1.35 1.15 3.35 1.50
Ours (Loose Coupling) 0.99 0.93 0.68 1.90 1.32 1.33 1.15 2.64 1.37
@@
Our evaluation encompasses three datasets: the NeRF synthetic dataset (Mildenhall et al., 2021), the DTU real-captured dataset (Jensen et al., 2014), and the Tanks & Temples (TNT) large-scale dataset (Knapitsch et al., 2017). The NeRF synthetic dataset consists of eight objects, each rendered via path tracing to generate 100 views. The DTU and TNT datasets are processed to obtain camera poses in accordance with (Gao et al., 2023), while the images and masks are processed following the methodology outlined in (Liu et al., 2020a). The image resolutions for the DTU and TNT datasets are $800\times 600$ and $1920\times 1080$, respectively.
Comparisons
We benchmark our method on two tasks: surface reconstruction and novel view synthesis. Our approach is compared against several state-of-the-art 1) Neural surface reconstruction methods, including NeuS (Wang et al., 2021), NeuSG (Chen et al., 2023), RelightableGaussian (Gao et al., 2023), NeRF2Mesh (Tang et al., 2023b), Neuralangelo (Li et al., 2023b), NeUS2 (Wang et al., 2023), BakedSDF (Yariv et al., 2023), VolSDF (Yariv et al., 2021), and 2) Gaussian-based reconstruction methods such as SuGaR (Guédon and Lepetit, 2023), 2DGS (Huang et al., 2024), and GOF (Yu et al., 2024b). It is worth noting that GOF (Yu et al., 2024b) is a concurrent work. Additionally, we also compare the novel view synthesis capabilities of our method with works that focus solely on novel view synthesis tasks, including NeRF (Mildenhall et al., 2021), Mip-NeRF (Barron et al., 2021), Ins-NGP (Müller et al., 2022), NeUS2 (Wang et al., 2023), and 3DGS (Kerbl et al., 2023). The quality of image synthesis is quantified using PSNR. For surface reconstruction accuracy, we employ Chamfer distance and the F1 score (Knapitsch et al., 2017).
@@
SuGaR (Guédon and Lepetit, 2023) 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33
2DGS (Huang et al., 2024) 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80
GOF (Yu et al., 2024b) (concurrent) 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74
Ours 0.44 0.96 0.40 0.36 1.02 0.80 0.64 1.20 1.08 0.97 0.54 0.72 0.37 0.52 0.42 0.70
@@
Table 2. Quantitative assessment on the DTU dataset with each cell colored to indicate the best second best and third best. Our method achieves the highest average quality of surface reconstruction and the lowest Chamfer-$\mathcal{L}_{1}$ distance with real-world data.
@@
Figure 6. Qualitative comparisons of reconstructed results between Relightable 3D Gaussian (Gao et al., 2023), 2DGS (Huang et al., 2024) and our method on the DTU dataset. Our method yields the highest-quality surface reconstruction preserving fine details and intricate structures.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
NeRF (Barron et al., 2023) 31.74 0.947 0.056
NeUS2 (Wang et al., 2023) 30.39 0.944 0.064
3D-GS (Kerbl et al., 2023) 33.32 0.970 0.031
2DGS (Huang et al., 2024) 33.07 0.967 0.033
Ours 33.86 0.970 0.031
Table 3. More quantitative assessment on the NeRF synthetic dataset.
5.2. Main Results and Comparisons
Quantitative analysis
Tables 12 and 3 present the quantitative results for NeRF-synthetic and DTU datasets, respectively. On the NeRF-synthetic dataset (Table 1 and Table 3), our approach achieves the best performance in both 3D surface reconstruction and novel view synthesis. Notably, our approach even outperforms the leading neural surface reconstruction methods for reconstruction evaluation while being much faster in training and yielding much better novel view synthesis results. Besides, our approach also delivers higher novel-view synthesis results than 3DGS (Kerbl et al., 2023) and 2DGS (Huang et al., 2024) in terms of PSNR, while achieving performance comparable to 3DGS on SSIM and LPIPS. On the DTU dataset (Table 2), our method also exhibits the best performance in terms of average surface reconstruction quality when compared with Gaussian-based reconstruction methods. It also performs favorably when compared with leading neural implicit reconstruction methods, which take much longer time to train and often yield inferior rendering qualities. We find that existing methods that achieve better reconstruction qualities tend to sacrifice the rendering qualities with lower performance than the NeRF baseline. This might be attributed to the different levels of detail required by surface reconstruction and novel view synthesis. Overall, our 3DGSR achieves excellent performance in both surface reconstruction and novel view synthesis.
Qualitative analysis
We provide qualitative comparisons of the reconstruction results in Figures 6 and 8. The state-of-the-art NeRF-based 3D reconstruction method, NeuS (Wang et al., 2021), fails to preserve fine details such as intricate Mic textures, Lego tracks, and stripes on the Hotdog disk. The 2DGS method (Huang et al., 2024), which utilizes 2D Gaussians and relies on an additional post-processing depth fusion step to obtain 3D surfaces, tends to produce over-smoothed reconstruction results, as evidenced by the smoothed Lego track and house. GOF (Yu et al., 2024b) employs Gaussian opacities and volumetric rendering to extract meshes directly from 3DGS, unifying geometry and appearance modeling for 3DGS. However, due to the inherent differences in the level of detail for geometry and appearance modeling, their results tend to produce noisy and non-existent structures needed for modeling high-frequency appearances, as seen in the noisy balls and Lego tracks in Figure 8. In contrast, our approach is capable of producing high-quality reconstructions, effectively preserving the finer details and maintaining the overall structure of the objects. More qualitative results and comparisons are included in Figure 9 and 10.
Speed comparison
We compared the training and inference speed of our method with Neuralangelo (Li et al., 2023b) and 2DGS (Huang et al., 2024), as shown in Table 4. When training on the DTU dataset, neural surface reconstruction methods like Neuralangelo (Li et al., 2023b) require over 30 hours to produce the final results, while Gaussian-splatting-based methods like 2DGS (Huang et al., 2024) take only 18.8 minutes. Our method, 3DGSR, takes 30 minutes-slightly longer than 2DGS (Huang et al., 2024), but significantly faster than Neuralangelo (Li et al., 2023b). This discrepancy arises because 3DGSR incorporates geometric constraints that effectively optimize the implicit representation, though querying the rays during training incurs additional time costs. However, when rendering an 800x600 image on the DTU dataset, both 2DGS (Huang et al., 2024) and our method achieve a frame rate of 175 fps. This is because surface regularization is applied solely during the training phase, allowing for the simultaneous learning of high-quality surfaces and Gaussians. Once the optimization is complete, the surface can be extracted offline and reused for future applications.
Method Training Speed Inference
Neuralangelo (Li et al., 2023b) ¿ 30 h ¡ 1 FPS
2DGS (Huang et al., 2024) 18.8 min 175 FPS
Ours 30 min 175 FPS
Table 4. Training and inference speed comparison on DTU dataset.
Method C-$\mathcal{L}_{1}$ PSNR LPIPS SSIM
static $\beta$ (Tight coupling) 1.95 33.02 0.0385 0.966
Poisson (Tight coupling) 2.03 - - -
Ours (Tight coupling) 1.50 33.23 0.0340 0.968
Ours (Loose coupling) 1.37 33.86 0.0330 0.971
w/o depth,normal 3.63 32.83 0.0380 0.963
w/o normal 1.74 33.18 0.0383 0.967
Table 5. Ablation study of the surface and GS alignment on the NeRF Synthetic dataset and the volumetric rendering depth (normal) loss, where we evaluate the effect of each proposed component. C-$\mathcal{L}_{1}$ is an abbreviation for the Chamfer-$\mathcal{L}_{1}$ distance.
5.3. Ablation Studies